Show that the sequence (an ),where an [ n/(n2+4) ] where is monotonic. Is (an ) a Cauchy sequence? Justify your answer.


1
Expert's answer
2021-06-15T09:05:47-0400

"\displaystyle a_n = \frac{n}{n^2+4}"

The sequence is monotonic if "a_{n+1} < a_n \Rightarrow (a_{n+1}-a_n) <0".


"\displaystyle a_{n+1}=\frac{n+1}{(n+1)^2+4}=\frac{n+1}{n^2 +2n+5}"


"\displaystyle \frac{n+1}{n^2+2n+5} - \frac{n}{n^2+4} \\[9pt]= \frac{(n+1)(n^2+4)-(n^2+2n+5)n}{(n^2+2n+5)(n^2+4)}\\[9pt] = \frac{n^3+4n+n^2+4 -n^3-2n^2-5n}{(n^2+2n+5)(n^2+4)}\\[9pt] = \frac{-n^2-n+4}{(n^2+2n+5)(n^2+4)}"

"\\[9pt]-(n^2+n-4) <0" is true for any "n\geq 2".


Nevertheless, this sequence is not monotonic because "a_2>a_1".


This sequence is Cauchy sequence because there exists such a "n=n_0" that for all "n>n_0, p>n_0"


 and any "\epsilon>0:" "| a_{n+p} -a_n| < \epsilon" . Starting from n=2 this indeed works for any "\epsilon>0".